OpenStudy (aravindg):

physics help!!!!!!!1

OpenStudy (aravindg):

i need help n motion

OpenStudy (anonymous):

?

OpenStudy (aravindg):

problem on pulley and friction

OpenStudy (anonymous):

ok

OpenStudy (aravindg):

can u help?

OpenStudy (anonymous):

yes

OpenStudy (aravindg):

A particle of mass 1gm executes an oscillatory motion on the concave surface of spherical dish of radius 2 m placed on a horizontal plane.If the motion of the particle begins from a point on the dishat a height of 1cm from the horizontal plane and the coefficient of friction is 0.01 ,find the total distance covered by the particle before it comes to rest.

OpenStudy (anonymous):

i might be wrong but all the potential energy turn to the work of the friction force mgh = rtheta*frictionforce mgh = rtheta*mg*0.01 h/(0.01r) = theta 1/2 = theta so the distance covered 1/2 * r = 1m

OpenStudy (anonymous):

tell me if you agree cause i might be wrong.

OpenStudy (anonymous):

actually N is mgcostheta

OpenStudy (anonymous):

but it almost always 1 .. you can solve with costheta but the answer wont dfffer much